# CoReEcho: Continuous Representation Learning for 2D+time Echocardiography Analysis

### v1

@@ Abstract
Deep learning (DL) models have been advancing automatic medical image analysis on various modalities, including echocardiography, by offering a comprehensive end-to-end training pipeline. This approach enables DL models to regress ejection fraction (EF) directly from 2D+time echocardiograms, resulting in superior performance. However, the end-to-end training pipeline makes the learned representations less explainable. The representations may also fail to capture the continuous relation among echocardiogram clips, indicating the existence of spurious correlations, which can negatively affect the generalization. To mitigate this issue, we propose CoReEcho, a novel training framework emphasizing continuous representations tailored for direct EF regression. Our extensive experiments demonstrate that CoReEcho: 1) outperforms the current state-of-the-art (SOTA) on the largest echocardiography dataset (EchoNet-Dynamic) with MAE of 3.90 & R2 of 82.44, and 2) provides robust and generalizable features that transfer more effectively in related downstream tasks. The code is publicly available at https://github.com/fadamsyah/CoReEcho.
@@ 3 Experiments
We conducted experiments using three standard echocardiography benchmarks, namely EchoNet-Dynamic [18], CAMUS [10], and HMC-QU [3]. We tested the CoReEcho training framework on [18] to demonstrate its efficacy and then performed transfer learning experiments on [10] and [3] to investigate the CoReEcho pretrained model transferability. [18] and [10] primarily deal with EF regression, whereas [3] focuses on MI classification. The preprocessing, augmentation, and hyperparameter details, are provided in the Appendix.
@@ 4.1 Ejection Fraction Regression
Table 1: EchoNet-Dynamic: Comparison with SOTA methods on the test set. CoReEcho achieves the best performance (${R2=82.43}$). ($*{}^{*}$): reproduced performance. SH: single heartbeat
@@
$\mathcal{L}_{1st}$
$\mathcal{A}$
$2\times\mathcal{C}$
R2
@@
Ours Tuning CoReEcho 0.807 5.29 0.693 6.81
@@
Ours Fine- Tuning CoReEcho 92.95 86.07 93.55 92.91 90.64
@@
Comparison with SOTA. In Table 5, we compare CoReEcho with SOTA methods using the largest echocardiography dataset [18]. While assessing both CoReEcho and EchoCoTr [16], we average the EF predictions of three sampled clips for each video to reduce the randomness introduced by the clip sampling process. CoReEcho achieves the lowest MAE & RMSE, and the highest R2, indicating superiority over other methods. Additionally, CoReEcho is more efficient than [18, 19, 25] since it provides a direct EF regression, thus removing the need for the segmentation stage to determine a heartbeat cycle.
@@ 5 Conclusion
We propose CoReEcho, a novel training framework for a single-view echocardiography video EF estimation. CoReEcho enables the encoder to represent the continuity relation across samples in the embedding space and train a shallow MLP head to leverage the extracted features to predict EF. CoReEcho outperforms the state-of-the-art methods on the EchoNet-Dynamic test set. Our experiments further show that CoReEcho can provide a strong pretrained model which extracts robust and universal features for single-view echocardiography downstream tasks. These observations indicate that allowing a DL-based echocardiography diagnosis model to learn more explainable representation elevates the generalization of in-domain and OOD data. Based on this, we believe CoReEcho can serve as an intermediate fine-tuning task for developing an echocardiography foundation model. An intermediate fine-tuning task is critical to tailor foundation model features to be more suitable for related downstream tasks, as evidenced by several works conducted with natural [26] and medical data [1].
@@
Table 7: Train hyperparameters. ($*{}^{*}$): 25 epochs for the 1st training stage and 5 epochs for the 2nd stage.
@@
Table 7: Train hyperparameters. ($*{}^{*}$): 25 epochs for the 1st training stage and 5 epochs for the 2nd stage.
Figure 4: Grad-CAM. CoReEcho $\mathcal{F}_{E}$ exhibits reduced focus on backgrounds and more focus on the LV.

### v2

@@ Abstract
Deep learning (DL) models have been advancing automatic medical image analysis on various modalities, including echocardiography, by offering a comprehensive end-to-end training pipeline. This approach enables DL models to regress ejection fraction (EF) directly from 2D+time echocardiograms, resulting in superior performance. However, the end-to-end training pipeline makes the learned representations less explainable. The representations may also fail to capture the continuous relation among echocardiogram clips, indicating the existence of spurious correlations, which can negatively affect the generalization. To mitigate this issue, we propose CoReEcho, a novel training framework emphasizing continuous representations tailored for direct EF regression. Our extensive experiments demonstrate that CoReEcho: 1) outperforms the current state-of-the-art (SOTA) on the largest echocardiography dataset (EchoNet-Dynamic) with MAE of 3.90 & R2 of 82.44, and 2) provides robust and generalizable features that transfer more effectively in related downstream tasks. The code is publicly available at https://github.com/BioMedIA-MBZUAI/CoReEcho.
@@ 3 Experiments
We used an RTX 4090 (24 GB) with PyTorch 1.12.1 and CUDA 11.6 throughout our experiments. We conducted experiments using three standard echocardiography benchmarks, namely EchoNet-Dynamic [19], CAMUS [11], and HMC-QU [4]. We tested the CoReEcho training framework on [19] to demonstrate its efficacy and then performed transfer learning experiments on [11] and [4] to investigate the CoReEcho pretrained model transferability. [19] and [11] primarily deal with EF regression, whereas [4] focuses on MI classification. The preprocessing, augmentation, and hyperparameter details, are provided in the Appendix.
@@ 4.1 Ejection Fraction Regression
Table 1: EchoNet-Dynamic: Comparison with SOTA methods on the test set. CoReEcho achieves the best performance. (): reproduced performance. SH: single heartbeat
@@
$\mathcal{L}_{1st}$ $\mathcal{A}$ $2\times\mathcal{C}$ R2
@@
Our Implementation Tuning CoReEcho 0.807 5.29 0.693 6.81
@@
Our Implementation Fine- Tuning CoReEcho 92.95 86.07 93.55 92.91 90.64
@@
Comparison with SOTA. In Table 5, we compare CoReEcho with SOTA methods using the largest echocardiography dataset [19]. While assessing both CoReEcho and EchoCoTr [17], we average the EF predictions of three sampled clips for each video to reduce the randomness introduced by the clip sampling process. CoReEcho achieves the lowest MAE of 3.90 and RMSE of 5.13, and the highest R2, indicating superiority over other methods. Additionally, CoReEcho is more efficient than [19, 20, 26] since it provides a direct EF regression, thus removing the need for the segmentation stage to determine a heartbeat cycle. The training time for CoReEcho is 4 hours, and the inference time is 5.95 $\pm$ 0.14 ms per clip, demonstrating our method’s suitability for real-time processing.
@@ 5 Conclusion
We propose CoReEcho, a novel training framework for a single-view echocardiography video EF estimation. CoReEcho enables the encoder to represent the continuity relation across samples in the embedding space and train a shallow MLP head to leverage the extracted features to predict EF. CoReEcho outperforms the state-of-the-art methods on the EchoNet-Dynamic test set. Our experiments further show that CoReEcho can provide a strong pretrained model which extracts robust and universal features for single-view echocardiography downstream tasks. These observations indicate that allowing a DL-based echocardiography diagnosis model to learn more explainable representation elevates the generalization of in-domain and OOD data. Based on this, we believe CoReEcho can serve as an intermediate fine-tuning task for developing an echocardiography foundation model. An intermediate fine-tuning task is critical to tailor foundation model features to be more suitable for related downstream tasks, as evidenced by several works conducted with natural [27] and medical data [1]. In addition, this continuity in representations may be used to detect distribution shifts during deployment, enabling subsequent domain adaptation to maintain performance. Exploring these topics could enhance real-world diagnostic processes, which we leave for future work.
{credits}
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
@@
Table 7: Train hyperparameters. (): 25 epochs for the 1st training stage and 5 epochs for the 2nd stage.
@@
[…]
figure Grad-CAM. CoReEcho $\mathcal{F}_{E}$ exhibits reduced focus on backgrounds and more focus on the LV.